OpenStudy (anonymous):

Solve for x: x/a-x-a-x/x=a/x . . . ANSWER {2a/3}

OpenStudy (anonymous):

\[\frac{ x }{ a-x }-\frac{ a-x }{ x }=\frac{ a }{ x }\]

OpenStudy (zzr0ck3r):

multiply everything by x*(a-x) and tell me what oyu get

OpenStudy (zzr0ck3r):

you get*

OpenStudy (shubhamsrg):

=> (x-a+a)/(a-x) - (a/x - x/x) = a/x =>( (x-a)/(a-x) ) + a/(a-x) - a/x -1 =a/x =>-1 + a/(a-x) -1 = 2a/x =>a/(a-x) = 2a/x + 2 easier to cross multiply now..though cross multiplying in the initial most condition wouldnt have made much difference either..just would have been lengthy..

OpenStudy (anonymous):

x/(a-x) -(a/x -x/x)=a/x or x/(a-x) -a/x+1=a/x or(x+a-x)/(a-x) = a/x+a/x or a/(a-x)=2a/x or 2(a-x)=x or 2a-2x=x or 2a=3x or x=2a/3